• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E.Chudakov

    Log entry time 20:32:14 on November 19, 2008

    Entry number 249793

    keyword=Experiment: summary for 11/18(day)-11/19(day)

    The experiment is taking data for NegNeg polarity, vertical target
    polarization.
    The target polarization is more than 60%, the beam current is 12uA.
    The data look OK.

    Beam time accounting in hours:
    date ABU BANU BNA losses
    2008/11/18 day 2.32 3.59 2.08 unclear
    2008/11/18 swing 0.00 6.00 2.00 control access 6h - Q2R(short), fire
    alarm (long)
    2008/11/19 owl 6.62 0.42 0.95 smooth running
    2008/11/19 day 4.05 0.99 2.94 Q1 issues

    Technical problems:
    11/18 16:00 - a compressor for the dump dehumidifier started smoking.
    Dave Kausch turned it off. It has to be repaired soon.
    11/18 18:00 - a problem with the Q2R helium meter was fixed.
    11/19 day - Q1 trips more and more frequently

    Status of commissioning of other equipment:
    Gas Cherenkov: the high background in the PMTs is being investigated
    RICH: DAQ problems
    Compton electron detector: DAQ problems


    A copy of this log entry has been emailed to: transversity@jlab.org